Running from the top flagpole to a hook in the ground there is a rope that is 10 meters long. If the hook is 8 meters from the base of the flagpole, how tall is the flagpole?


1
Expert's answer
2021-12-07T09:08:32-0500

Height of the flag pole ;

"10^2 =8^2 +H^2"

"H^2 =100-64"

"H^2 =36"

"H=\sqrt{36}"

"H=6m"


Height of the flag pole =6m